AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 11 of 33  ELOG logo
  ID Date Author Subject Text Attachments
  465   Sat Jan 13 14:25:05 2024 CB, JMBias tests with all FEEs connectedConnected all FEEs.

The two top fees corresponding
to the bottom right detector are connected
 Screenshot_2024-01-13_15-24-40.png20240113_150158.jpg 
  464   Thu Jan 11 19:57:20 2024 CB, JMAll FEE holders mountedAfter discussions with Davide, changed
the holes in the FEE holders for the bottom-right
detector to slots in order to slightly change
  
  463   Wed Jan 10 19:27:08 2024 CB, JMManifold and more bias testsMounted water manifold on CARME with help
from GSI plumber. Connected all the hoses
with Swagelok self-locking connectors. Tested
  
  462   Wed Jan 10 10:22:41 2024 CB, JMSome FEEs mounted - issues with Det#2 (bottom left)Started mounting some FEEs. We ran into
some mechanical issues due to the new ERNI
connectors on the new flexi-rigid adaptor
 Screenshot_2024-01-10_11-59-39.pngScreenshot_2024-01-10_12-56-36.png 
  461   Mon Jan 8 19:29:57 2024 CB, JMManifold repaired + fee assembly in bench Davide extracted the stuck blind plugs
and re-threaded all missing ports. Completed
manifold assembly, except for main inlet/outlet
 20240108_164756.jpg20240108_163218.jpg20240108_163916.jpg20240108_164010.jpg 
  460   Mon Jan 8 19:27:36 2024 CB, JMLeak current tests overnight Left bias in overnight. Leakage current
decreased and stabilised. See attached. Bias
off as we prepare to mount the fees.
 Screenshot_2024-01-08_09-56-20.png 
  459   Sun Jan 7 15:38:57 2024 CB, JMManifolds mostly assembled + bias overnight checkAssembled most of the connectors on the
new (green) manifolds. Two blanking pulgs
stuck when mounting - cannot remove them
 Screenshot_2024-01-07_16-38-26.png 
  458   Fri Jan 5 12:29:10 2024 JM, CBFurther bias testsBiased detectors this morning to 150V,
the leakage current for all detectors appears
stable and is ~3.5 uA. We have also replaced
  
  457   Thu Jan 4 13:15:10 2024 CB, JMTent removed - bias testRemoved tent, octagon, Neg cables.

Disconnected and removed Baking
Pi, TC08, thermocouples.
 Bias.png 
  456   Wed Oct 18 16:47:36 2023 JG, CBPower outageLocal power outage at the CRYRING today,
while JG was trying to activate the ion pumps.
Unlcear what caused it.
  
  455   Tue Oct 17 08:43:47 2023 CBNEG activation -209:40 - Found temperature and pressure
stablised, but too high to comfortably activate
NEGs closest to detectors. Detector temperature
 Screenshot_2023-10-17_09-45-23.pngScreenshot_2023-10-17_11-04-09.pngScreenshot_2023-10-17_14-56-55.pngScreenshot_2023-10-17_17-58-24.pngScreenshot_2023-10-17_17-58-51.png 
  454   Mon Oct 16 08:28:02 2023 CBNEG activation -19:00 - Michael's online vacuum pressure
plot is no longer working - an old image
is being displayed. CARME temperatures around
 Screenshot_2023-10-16_17-17-30.png 
  453   Sun Oct 15 09:35:03 2023 CBStarting ramp down10:40 - Status as attached. Heaters setpoint/read
at 250 C, 150 C, 225 C. Power stable at 53%,
41%, 63%
 Screenshot_2023-10-15_10-35-47.pngScreenshot_2023-10-15_17-56-00.png 
  452   Tue Oct 3 08:48:36 2023 TDTuesday 3 October09.40 CARME IE514 1.79e-8 mbar, ambient temperature
20.7 deg C
 1000007432.jpg1000007431.jpg1000007434.jpg 
  451   Mon Oct 2 08:41:43 2023 TDMonday 2 October09.35 CARME IE514 1.95e-8 mbar, ambient temperature
20.8 deg C
 6x 
  450   Sun Oct 1 10:24:50 2023 TDSunday 1 October11.20 CARME IE514 3.53e-9 mbar, ambient temperature
20.6 deg C
 12x 
  449   Sat Sep 30 10:58:14 2023 TDSaturday 30 September10.50 CARME IE514 7.14e-10 mbar, ambient temperature
20.8 deg C
 9x 
  448   Fri Sep 29 08:54:29 2023 TDFriday 29 September
09.50 CARME IE514 9.0e-11 mbar, ambient temperature
21.3 deg C
 11x 
  447   Thu Sep 28 09:19:50 2023 TDThursday 28 September
10.20 CARME IE514 8.6e-11 mbar
 1000007142.jpg1000007143.jpg 
  446   Wed Sep 27 15:56:41 2023 TDWednesday 27 September
16.55 CARME IE514 8.2e-11 mbar
 1000007129.jpg1000007127.jpg1000007128.jpg 
ELOG V3.1.4-unknown